Consider the formation of ammonia: N2(g) + 3 H2(g) 2NH3(g). (a). Assuming that H° and S° are constant with temperature, find the temperature at which Kp= 1.00.

(b). Find Kp at 400.0 °C, the typical temperature for NH3 production.


1
Expert's answer
2021-07-01T05:51:05-0400

Part a

"N_2(g)+ 3 H_2(g) = 2NH_3(g)"

By using values of enthalpy and entropy from table

"\u2206H\u00b0rxn=2\u00d7\u2206Hf(NH_3)-3\u00d7\u2206Hf(H_2)-\u2206Hf(N_2)"

"\u2206H\u00b0rxn=2\u00d7-45.9-3\u00d70-0"

∆H°rxn=-91.8 kJ/mol

"\u2206S\u00b0rxn=2\u00d7S\u00b0(NH3)-S\u00b0(N_2)-3\u00d7S\u00b0(H_2)"

"\u2206S\u00b0rxn=2\u00d7199-191.5-3\u00d7130.6"

"\u2206S\u00b0rxn=-197.3 J\/K"

By using relationship

"\u2206H\u00b0-T\u2206S\u00b0=-RTLnKp"

Kp=1

"-91.8\u00d710\u00b3+T\u00d7197.3=0"

"T=465.28 K"

Part b)

"At T=400\u00b0C=673 K"

∆H°-T∆S°=-RTLnKp

-91.8×10³-673×-197.3=-8.314×673×lnKp

"lnKp=-7.3245"

"K_p=6.592\u00d710^{-4}"